OpenStudy (anonymous):

proof that x/x+3=1-3/x+3 for each real number violates -3

OpenStudy (anonymous):

i've tried to factor but didn't got a solution :/ x(x+3)/x+3=x^2+3x/x+3 dunno what else to do x)

OpenStudy (radar):

On the right of the equal sign, is the -3 to be divided only by x or the quantity x+3?

OpenStudy (radar):

Same goes for the x on left side is it x/x+3 or x/(x+3)

OpenStudy (anonymous):

it's x/x+3 and 3/x+3 it's without the (1-) i mean it's like this 1-(3/x+3) so 3 is not negative

OpenStudy (radar):

If that is truly the case, then x/x is 1 so it would be this: 1+3 =1-3/x+3 4=4-3/x 0=-3/x This is a contradiction and I believe you are expressing it wrong or it is indeed some kind of violation.

OpenStudy (radar):

I guess you could multiply both sides by x and get 0=-3 which further proves the invalidity of this equation.

OpenStudy (anonymous):

i guess you didn't understand the question cuz what you have to do is find a way to make x/x+3 equal 1- 3/x+3 i mean they already gave you the result of x/x+3 which is 1- 3/x+3 and all what you have to do is to proof it :) imagine you don't have that result and you have to make x/x+3 give you that result :)

OpenStudy (radar):

You're right, I didn't understand the point. Nor do I understand it now lol.

OpenStudy (anonymous):

lol for real o.o